# Beyond Discretization: Learning the Optimal Solution Path

### v2

@@ A.1 Proofs from Section 2
By strong convexity of $h(\cdot,\lambda)$ and definition of $\boldsymbol{\beta}_{\rm sp}^{\ast}$,
$\|\boldsymbol{\Phi}(\lambda)\boldsymbol{\beta}_{\rm sp}^{\ast}-\boldsymbol{\theta}^{*}(\lambda)\|^{2}\ \leq\ \frac{2}{\mu}\sup_{\lambda\in\Lambda}\left\{h(\boldsymbol{\Phi}(\lambda)\boldsymbol{\beta}_{\rm sp}^{\ast},\lambda)-h(\boldsymbol{\theta}^{*}(\lambda),\lambda)\right\}\ =\ \frac{2}{\mu}\epsilon_{\rm sp}(\boldsymbol{\beta}_{\rm sp}^{\ast})\ =\ \frac{2}{\mu}\epsilon_{\rm sp}^{\ast},$
@@
Next, we bound $\|\boldsymbol{\beta}_{\mathrm{avg}}^{\ast}-\boldsymbol{\beta}_{\rm sp}^{\ast}\|^{2}$. Observe that due to the optimality of $\boldsymbol{\theta}^{*}(\lambda)$ for each $\lambda$,
$\mathbb{E}\left[h(\boldsymbol{\Phi}(\tilde{\lambda})\boldsymbol{\beta}_{\rm sp}^{\ast},\tilde{\lambda})\right]-\mathbb{E}\left[h(\boldsymbol{\Phi}(\tilde{\lambda})\boldsymbol{\beta}_{\mathrm{avg}}^{\ast},\tilde{\lambda})\right]\ \leq\ \mathbb{E}\left[h(\boldsymbol{\Phi}(\tilde{\lambda})\boldsymbol{\beta}_{\rm sp}^{\ast},\tilde{\lambda})-h(\boldsymbol{\theta}^{*}(\tilde{\lambda}),\tilde{\lambda})\right]\ \leq\ \epsilon_{\rm sp}^{\ast}.$
On the other hand, from the optimality of $\boldsymbol{\beta}_{\mathrm{avg}}^{\ast}$ and strong-convexity of $F(\cdot)$,
$F(\boldsymbol{\beta}_{\rm sp}^{\ast})-F(\boldsymbol{\beta}_{\mathrm{avg}}^{\ast})\ =\ \mathbb{E}\left[h(\boldsymbol{\Phi}(\tilde{\lambda})\boldsymbol{\beta}_{\rm sp}^{\ast},\tilde{\lambda})\right]-\mathbb{E}\left[h(\boldsymbol{\Phi}(\tilde{\lambda})\boldsymbol{\beta}_{\mathrm{avg}}^{\ast},\tilde{\lambda})\right]\ \geq\ \frac{c\mu}{2}\|\boldsymbol{\beta}_{\mathrm{avg}}^{\ast}-\boldsymbol{\beta}_{\rm sp}^{\ast}\|^{2}.$
Combining this inequality with the previous one shows $\frac{c\mu}{2}\|\boldsymbol{\beta}_{\mathrm{avg}}^{\ast}-\boldsymbol{\beta}_{\rm sp}^{\ast}\|^{2}\ \leq\ \epsilon_{\rm sp}^{\ast},$ which implies that
$\|\boldsymbol{\beta}_{\mathrm{avg}}^{\ast}-\boldsymbol{\beta}_{\rm sp}^{\ast}\|^{2}\ \leq\ \frac{2\epsilon_{\rm sp}^{\ast}}{c\mu}.$
Substitute the above into Equation 6. Take expectation over sample paths on both sides, we conclude that
@@ A.2 Proofs from Section 3
The proof follows the proof of Theorem 3.1 of Gower et al. (2019); we include it for completeness under the assumption of the RWGC Definition 3.1. Let $\eta:=\frac{\bar{\eta}}{\rho}$ denote the step-size. For $t\geq 0$, we have
$\|\boldsymbol{w}_{t+1}-\boldsymbol{w}^{*}\|^{2}$ $=\|\boldsymbol{w}_{t}-\boldsymbol{w}^{*}-\eta\nabla g(\boldsymbol{w}_{t},z_{t})\|^{2}$
$=\|\boldsymbol{w}_{t}-\boldsymbol{w}^{*}\|^{2}-2\eta\nabla g(\boldsymbol{w}_{t},z_{t})^{\top}(\boldsymbol{w}_{t}-\boldsymbol{w}^{*})+\eta^{2}\|\nabla g(\boldsymbol{w}_{t},z_{t})\|^{2}.$
@@
$\mathbb{E}\left[\|\boldsymbol{w}_{t+1}-\boldsymbol{w}^{*}\|^{2}\;\middle|\;\boldsymbol{w}_{t}\right]~{}$ $=~{}\|\boldsymbol{w}_{t}-\boldsymbol{w}^{*}\|^{2}-2\eta\nabla G(\boldsymbol{w}_{t})^{\top}(\boldsymbol{w}_{t}-\boldsymbol{w}^{*})+\eta^{2}\mathbb{E}\left[\|\nabla g(\boldsymbol{w}_{t},z_{t})\|^{2}\;\middle|\;\boldsymbol{w}_{t}\right]$
$\leq~{}\|\boldsymbol{w}_{t}-\boldsymbol{w}^{*}\|^{2}-2\eta\nabla G(\boldsymbol{w}_{t})^{\top}(\boldsymbol{w}_{t}-\boldsymbol{w}^{*})+\eta^{2}[2\rho(G(\boldsymbol{w}_{t})-G^{*})+\sigma^{2}]$
$\leq~{}\|\boldsymbol{w}_{t}-\boldsymbol{w}^{*}\|^{2}+2\eta\left(G^{*}-G(\boldsymbol{w}_{t})-\tfrac{\mu_{g}}{2}\|\boldsymbol{w}_{t}-\boldsymbol{w}^{*}\|^{2}\right)+\eta^{2}[2\rho(G(\boldsymbol{w}_{t})-G^{*})+\sigma^{2}]$
$=~{}\left(1-\eta\mu_{g}\right)\|\boldsymbol{w}_{t}-\boldsymbol{w}^{*}\|^{2}+2\left(\eta^{2}\rho-\eta\right)(G(\boldsymbol{w}_{t})-G^{*})+\eta^{2}\sigma^{2}$
$\leq~{}\left(1-\eta\mu_{g}\right)\|\boldsymbol{w}_{t}-\boldsymbol{w}^{*}\|^{2}+\eta^{2}\sigma^{2}$ (7)
@@
$\mathbb{E}\left[\|\boldsymbol{w}_{t}-\boldsymbol{w}^{*}\|^{2}\right]~{}$ $\leq~{}(1-\eta\mu_{g})^{t}\|\boldsymbol{w}_{0}-\boldsymbol{w}^{*}\|^{2}+\eta^{2}\sigma^{2}\sum_{k=0}^{t-1}(1-\eta\mu_{g})^{k}$
$\leq~{}(1-\eta\mu_{g})^{t}\|\boldsymbol{w}_{0}-\boldsymbol{w}^{*}\|^{2}+\frac{\eta^{2}\sigma^{2}}{1-(1-\eta\mu_{g})}$
$=~{}(1-\eta\mu_{g})^{t}\|\boldsymbol{w}_{0}-\boldsymbol{w}^{*}\|^{2}+\frac{\eta\sigma^{2}}{\mu_{g}},$
@@
$\boldsymbol{\bar{\beta}}(\lambda)\in\arg\min_{\boldsymbol{\beta}\in\mathbb{R}^{p}}f(\boldsymbol{\beta},\lambda)=\arg\min_{\boldsymbol{\beta}\in\mathbb{R}^{p}}h(\boldsymbol{\Phi}(\lambda)\boldsymbol{\beta},\lambda).$
@@
$f(\boldsymbol{\beta},\lambda)\geq f(\boldsymbol{\bar{\beta}}(\lambda),\lambda)+\nabla f(\boldsymbol{\bar{\beta}}(\lambda),\lambda)^{\top}(\boldsymbol{\beta}-\boldsymbol{\bar{\beta}}(\lambda))+\frac{1}{2CL}\|\nabla f(\boldsymbol{\beta},\lambda)-\nabla f(\boldsymbol{\bar{\beta}}(\lambda),\lambda)\|^{2}.$
Using $\nabla f(\boldsymbol{\bar{\beta}}(\lambda),\lambda)=0$ for all $\lambda$, rearranging and recalling the definitions of $\boldsymbol{\beta}_{\mathrm{avg}}^{\ast}$ (Section 2) and $\boldsymbol{\beta}_{\rm sp}^{\ast}$ (proof of Theorem 2.4) yields
$\frac{1}{2CL}\|\nabla f(\boldsymbol{\beta},\lambda)\|^{2}~{}$ $\leq~{}f(\boldsymbol{\beta},\lambda)-f(\boldsymbol{\bar{\beta}}(\lambda),\lambda)$
$\leq~{}f(\boldsymbol{\beta},\lambda)-h(\boldsymbol{\theta}^{*}(\lambda),\lambda)$
$=~{}(f(\boldsymbol{\beta},\lambda)-f(\boldsymbol{\beta}_{\mathrm{avg}}^{\ast},\lambda))+(f(\boldsymbol{\beta}_{\mathrm{avg}}^{\ast},\lambda)-f(\boldsymbol{\beta}_{\rm sp}^{\ast},\lambda))+(f(\boldsymbol{\beta}_{\rm sp}^{\ast},\lambda)-h(\boldsymbol{\theta}^{*}(\lambda),\lambda))$
$\leq~{}(f(\boldsymbol{\beta},\lambda)-f(\boldsymbol{\beta}_{\mathrm{avg}}^{\ast},\lambda))+(f(\boldsymbol{\beta}_{\mathrm{avg}}^{\ast},\lambda)-f(\boldsymbol{\beta}_{\rm sp}^{\ast},\lambda))+\epsilon_{\rm sp}^{\ast}$
where the second inequality uses optimality of $\boldsymbol{\theta}^{*}(\lambda)$ and the third uses that $\epsilon_{\rm sp}^{\ast}$ is attained at $\boldsymbol{\beta}_{\rm sp}^{\ast}$. Taking expectations on both sides, we get
$\frac{1}{2CL}\mathbb{E}\left[\|\nabla f(\boldsymbol{\beta},\tilde{\lambda})\|^{2}\right]\ \leq\ F(\boldsymbol{\beta})-F(\boldsymbol{\beta}_{\mathrm{avg}}^{\ast})+F(\boldsymbol{\beta}_{\mathrm{avg}}^{\ast})-F(\boldsymbol{\beta}_{\rm sp}^{\ast})+\epsilon_{\rm sp}^{\ast}\ \leq\ F(\boldsymbol{\beta})-F(\boldsymbol{\beta}_{\mathrm{avg}}^{\ast})+\epsilon_{\rm sp}^{\ast},$
where we used that the definition of $\boldsymbol{\beta}_{\mathrm{avg}}^{\ast}$ implies that $F(\boldsymbol{\beta}_{\mathrm{avg}}^{\ast})\leq F(\boldsymbol{\beta}_{\rm sp}^{\ast})$. Multiplying through by $2CL$ completes the proof. ∎
@@
Our strategy will be to invoke Lemma 3.3 and apply Theorem 3.2, in conjuction with Theorem 2.4. To that end, notice that the upper bound on the step-size parameter $\bar{\eta}$ given in Theorem 3.2 becomes $\frac{CL}{c\mu}\geq 1$. Hence, we need only constrain $0<\bar{\eta}\leq 1$.
@@
$\mathbb{E}\left[\|\boldsymbol{\beta}_{T}-\boldsymbol{\beta}_{\mathrm{avg}}^{\ast}\|^{2}\right]\ \leq\ \left(1-\bar{\eta}\frac{c\mu}{CL}\right)^{T}\|\boldsymbol{\beta}_{0}-\boldsymbol{\beta}_{\mathrm{avg}}^{\ast}\|^{2}+\frac{2\bar{\eta}\epsilon_{\rm sp}^{\ast}}{c\mu}.$
Taking expectation of Theorem 2.4 on both sides and substituting the above, we have
$\mathbb{E}\left[\epsilon_{\rm sp}(\boldsymbol{\beta}_{T})\right]$ $\leq\ 2CL\mathbb{E}\left[\|\boldsymbol{\beta}_{T}-\boldsymbol{\beta}_{\mathrm{avg}}^{\ast}\|^{2}\right]+\frac{8CL}{c\mu}\epsilon_{\rm sp}^{\ast}$
$\leq\ 2CL\left(1-\bar{\eta}\frac{c\mu}{CL}\right)^{T}\|\boldsymbol{\beta}_{0}-\boldsymbol{\beta}_{\mathrm{avg}}^{\ast}\|^{2}+4\frac{\bar{\eta}CL}{c\mu}\epsilon_{\rm sp}^{\ast}+\frac{8CL}{c\mu}\epsilon_{\rm sp}^{\ast}.$ (8)
Rearranging yields the first result. We next prove the two bounds on the iteration complexity.
First consider the case $\epsilon_{\rm sp}^{\ast}>0$. Our goal will be to choose $T$ large enough to drive the first term on the right side of (8) to below $4\frac{CL}{c\mu}\bar{\eta}\epsilon_{\rm sp}^{\ast}$. To simplify exposition, let $\kappa=\frac{CL}{c\mu}$. Then, we need to ensure that
$T\log(1-\bar{\eta}/\kappa)\leq\log\left(\frac{2\kappa\bar{\eta}\epsilon_{\rm sp}^{\ast}}{CL\|\boldsymbol{\beta}_{0}-\boldsymbol{\beta}_{\mathrm{avg}}^{\ast}\|^{2}}\right).$
@@
$T\geq\frac{\kappa}{\bar{\eta}}\log\left(\frac{CL\|\boldsymbol{\beta}_{0}-\boldsymbol{\beta}_{\mathrm{avg}}^{\ast}\|^{2}}{2\kappa\bar{\eta}\epsilon_{\rm sp}^{\ast}}\right).$
@@
$\mathbb{E}\left[\epsilon_{\rm sp}(\boldsymbol{\beta}_{T})\right]$ $\leq\ 2CL\left(1-\bar{\eta}\frac{c\mu}{CL}\right)^{T}\|\boldsymbol{\beta}_{0}-\boldsymbol{\beta}_{\mathrm{avg}}^{\ast}\|^{2}.$
@@
$T\log(1-\bar{\eta}/\kappa)\leq\log\left(\frac{\epsilon}{2CL\|\boldsymbol{\beta}_{0}-\boldsymbol{\beta}_{\mathrm{avg}}^{\ast}\|^{2}}\right).$
@@
$T\geq\frac{\kappa}{\bar{\eta}}\log\left(\frac{2CL\|\boldsymbol{\beta}_{0}-\boldsymbol{\beta}_{\mathrm{avg}}^{\ast}\|^{2}}{\epsilon}\right).$

### v3

@@ A.1 Proofs from Section 2
By strong convexity of $h(\cdot,\lambda)$ and definition of $\bm{\beta}_{\rm sp}^{\ast}$,
$\|\bm{\Phi}(\lambda)\bm{\beta}_{\rm sp}^{\ast}-\bm{\theta}^{*}(\lambda)\|^{2}\ \leq\ \frac{2}{\mu}\sup_{\lambda\in\Lambda}\left\{h(\bm{\Phi}(\lambda)\bm{\beta}_{\rm sp}^{\ast},\lambda)-h(\bm{\theta}^{*}(\lambda),\lambda)\right\}\ =\ \frac{2}{\mu}\epsilon_{\rm sp}(\bm{\beta}_{\rm sp}^{\ast})\ =\ \frac{2}{\mu}\epsilon_{\rm sp}^{\ast},$
@@
Next, we bound $\|\bm{\beta}_{\mathrm{avg}}^{\ast}-\bm{\beta}_{\rm sp}^{\ast}\|^{2}$. Observe that due to the optimality of $\bm{\theta}^{*}(\lambda)$ for each $\lambda$,
$\mathbb{E}\left[h(\bm{\Phi}(\tilde{\lambda})\bm{\beta}_{\rm sp}^{\ast},\tilde{\lambda})\right]-\mathbb{E}\left[h(\bm{\Phi}(\tilde{\lambda})\bm{\beta}_{\mathrm{avg}}^{\ast},\tilde{\lambda})\right]\ \leq\ \mathbb{E}\left[h(\bm{\Phi}(\tilde{\lambda})\bm{\beta}_{\rm sp}^{\ast},\tilde{\lambda})-h(\bm{\theta}^{*}(\tilde{\lambda}),\tilde{\lambda})\right]\ \leq\ \epsilon_{\rm sp}^{\ast}.$
On the other hand, from the optimality of $\bm{\beta}_{\mathrm{avg}}^{\ast}$ and strong-convexity of $F(\cdot)$,
$F(\bm{\beta}_{\rm sp}^{\ast})-F(\bm{\beta}_{\mathrm{avg}}^{\ast})\ =\ \mathbb{E}\left[h(\bm{\Phi}(\tilde{\lambda})\bm{\beta}_{\rm sp}^{\ast},\tilde{\lambda})\right]-\mathbb{E}\left[h(\bm{\Phi}(\tilde{\lambda})\bm{\beta}_{\mathrm{avg}}^{\ast},\tilde{\lambda})\right]\ \geq\ \frac{c\mu}{2}\|\bm{\beta}_{\mathrm{avg}}^{\ast}-\bm{\beta}_{\rm sp}^{\ast}\|^{2}.$
Combining this inequality with the previous one shows $\frac{c\mu}{2}\|\bm{\beta}_{\mathrm{avg}}^{\ast}-\bm{\beta}_{\rm sp}^{\ast}\|^{2}\ \leq\ \epsilon_{\rm sp}^{\ast},$ which implies that
$\|\bm{\beta}_{\mathrm{avg}}^{\ast}-\bm{\beta}_{\rm sp}^{\ast}\|^{2}\ \leq\ \frac{2\epsilon_{\rm sp}^{\ast}}{c\mu}.$
Substitute the above into Equation 6. Take expectations over sample paths on both sides, we conclude that
@@ A.2 Proofs from Section 3
The proof follows the proof of Theorem 3.1 of Gower et al. (2019); we include it for completeness under the assumption of the RWGC Definition 3.1. Let $\eta:=\frac{\bar{\eta}}{\rho}$ denote the step-size, $\bm{w}^{*}:=\arg\min_{\bm{w}\in\mathbb{R}^{d}}G(\bm{w})$ denote the optimality input of $G$. For $t\geq 0$, we have
$\|\bm{w}_{t+1}-\bm{w}^{*}\|^{2}$ $=\|\bm{w}_{t}-\bm{w}^{*}-\eta\nabla g(\bm{w}_{t},z_{t})\|^{2}$
$=\|\bm{w}_{t}-\bm{w}^{*}\|^{2}-2\eta\nabla g(\bm{w}_{t},z_{t})^{\top}(\bm{w}_{t}-\bm{w}^{*})+\eta^{2}\|\nabla g(\bm{w}_{t},z_{t})\|^{2}.$
@@
$\mathbb{E}\left[\|\bm{w}_{t+1}-\bm{w}^{*}\|^{2}\;\middle|\;\bm{w}_{t}\right]~{}$ $=~{}\|\bm{w}_{t}-\bm{w}^{*}\|^{2}-2\eta\nabla G(\bm{w}_{t})^{\top}(\bm{w}_{t}-\bm{w}^{*})+\eta^{2}\mathbb{E}\left[\|\nabla g(\bm{w}_{t},z_{t})\|^{2}\;\middle|\;\bm{w}_{t}\right]$
$\leq~{}\|\bm{w}_{t}-\bm{w}^{*}\|^{2}-2\eta\nabla G(\bm{w}_{t})^{\top}(\bm{w}_{t}-\bm{w}^{*})+\eta^{2}[2\rho(G(\bm{w}_{t})-G^{*})+\sigma^{2}]$
$\leq~{}\|\bm{w}_{t}-\bm{w}^{*}\|^{2}+2\eta\left(G^{*}-G(\bm{w}_{t})-\tfrac{\mu_{g}}{2}\|\bm{w}_{t}-\bm{w}^{*}\|^{2}\right)+\eta^{2}[2\rho(G(\bm{w}_{t})-G^{*})+\sigma^{2}]$
$=~{}\left(1-\eta\mu_{g}\right)\|\bm{w}_{t}-\bm{w}^{*}\|^{2}+2\left(\eta^{2}\rho-\eta\right)(G(\bm{w}_{t})-G^{*})+\eta^{2}\sigma^{2}$
$\leq~{}\left(1-\eta\mu_{g}\right)\|\bm{w}_{t}-\bm{w}^{*}\|^{2}+\eta^{2}\sigma^{2}$ (7)
@@
$\mathbb{E}\left[\|\bm{w}_{t}-\bm{w}^{*}\|^{2}\right]~{}$ $\leq~{}(1-\eta\mu_{g})^{t}\|\bm{w}_{0}-\bm{w}^{*}\|^{2}+\eta^{2}\sigma^{2}\sum_{k=0}^{t-1}(1-\eta\mu_{g})^{k}$
$\leq~{}(1-\eta\mu_{g})^{t}\|\bm{w}_{0}-\bm{w}^{*}\|^{2}+\frac{\eta^{2}\sigma^{2}}{1-(1-\eta\mu_{g})}$
$=~{}(1-\eta\mu_{g})^{t}\|\bm{w}_{0}-\bm{w}^{*}\|^{2}+\frac{\eta\sigma^{2}}{\mu_{g}},$
@@
$\bm{\bar{\beta}}(\lambda)\in\arg\min_{\bm{\beta}\in\mathbb{R}^{p}}f(\bm{\beta},\lambda)=\arg\min_{\bm{\beta}\in\mathbb{R}^{p}}h(\bm{\Phi}(\lambda)\bm{\beta},\lambda).$
@@
$f(\bm{\beta},\lambda)\geq f(\bm{\bar{\beta}}(\lambda),\lambda)+\nabla f(\bm{\bar{\beta}}(\lambda),\lambda)^{\top}(\bm{\beta}-\bm{\bar{\beta}}(\lambda))+\frac{1}{2CL}\|\nabla f(\bm{\beta},\lambda)-\nabla f(\bm{\bar{\beta}}(\lambda),\lambda)\|^{2}.$
Using $\nabla f(\bm{\bar{\beta}}(\lambda),\lambda)=0$ for all $\lambda$, rearranging and recalling the definitions of $\bm{\beta}_{\mathrm{avg}}^{\ast}$ (Section 2) and $\bm{\beta}_{\rm sp}^{\ast}$ (proof of Theorem 2.4) yields
$\frac{1}{2CL}\|\nabla f(\bm{\beta},\lambda)\|^{2}~{}$ $\leq~{}f(\bm{\beta},\lambda)-f(\bm{\bar{\beta}}(\lambda),\lambda)$
$\leq~{}f(\bm{\beta},\lambda)-h(\bm{\theta}^{*}(\lambda),\lambda)$
$=~{}(f(\bm{\beta},\lambda)-f(\bm{\beta}_{\mathrm{avg}}^{\ast},\lambda))+(f(\bm{\beta}_{\mathrm{avg}}^{\ast},\lambda)-f(\bm{\beta}_{\rm sp}^{\ast},\lambda))+(f(\bm{\beta}_{\rm sp}^{\ast},\lambda)-h(\bm{\theta}^{*}(\lambda),\lambda))$
$\leq~{}(f(\bm{\beta},\lambda)-f(\bm{\beta}_{\mathrm{avg}}^{\ast},\lambda))+(f(\bm{\beta}_{\mathrm{avg}}^{\ast},\lambda)-f(\bm{\beta}_{\rm sp}^{\ast},\lambda))+\epsilon_{\rm sp}^{\ast}$
where the second inequality uses optimality of $\bm{\theta}^{*}(\lambda)$ and the third uses that $\epsilon_{\rm sp}^{\ast}$ is attained at $\bm{\beta}_{\rm sp}^{\ast}$. Taking expectations on both sides, we get
$\frac{1}{2CL}\mathbb{E}\left[\|\nabla f(\bm{\beta},\tilde{\lambda})\|^{2}\right]\ \leq\ F(\bm{\beta})-F(\bm{\beta}_{\mathrm{avg}}^{\ast})+F(\bm{\beta}_{\mathrm{avg}}^{\ast})-F(\bm{\beta}_{\rm sp}^{\ast})+\epsilon_{\rm sp}^{\ast}\ \leq\ F(\bm{\beta})-F(\bm{\beta}_{\mathrm{avg}}^{\ast})+\epsilon_{\rm sp}^{\ast},$
where we used that the definition of $\bm{\beta}_{\mathrm{avg}}^{\ast}$ implies that $F(\bm{\beta}_{\mathrm{avg}}^{\ast})\leq F(\bm{\beta}_{\rm sp}^{\ast})$. Multiplying through by $2CL$ completes the proof. ∎
@@
Our strategy will be to invoke Lemma 3.3 and apply Theorem 3.2, in conjunction with Theorem 2.4. To that end, notice that the upper bound on the step-size parameter $\bar{\eta}$ given in Theorem 3.2 becomes $\frac{CL}{c\mu}\geq 1$. Hence, we need only constrain $0<\bar{\eta}\leq 1$.
@@
$\mathbb{E}\left[\|\bm{\beta}_{T}-\bm{\beta}_{\mathrm{avg}}^{\ast}\|^{2}\right]\ \leq\ \left(1-\bar{\eta}\frac{c\mu}{CL}\right)^{T}\|\bm{\beta}_{0}-\bm{\beta}_{\mathrm{avg}}^{\ast}\|^{2}+\frac{2\bar{\eta}\epsilon_{\rm sp}^{\ast}}{c\mu}.$
Taking the expectation of Theorem 2.4 on both sides and substituting the above, we have
$\mathbb{E}\left[\epsilon_{\rm sp}(\bm{\beta}_{T})\right]$ $\leq\ 2CL\mathbb{E}\left[\|\bm{\beta}_{T}-\bm{\beta}_{\mathrm{avg}}^{\ast}\|^{2}\right]+\frac{8CL}{c\mu}\epsilon_{\rm sp}^{\ast}$
$\leq\ 2CL\left(1-\bar{\eta}\frac{c\mu}{CL}\right)^{T}\|\bm{\beta}_{0}-\bm{\beta}_{\mathrm{avg}}^{\ast}\|^{2}+4\frac{\bar{\eta}CL}{c\mu}\epsilon_{\rm sp}^{\ast}+\frac{8CL}{c\mu}\epsilon_{\rm sp}^{\ast}.$ (8)
Rearranging yields the first result. Next, we prove the two bounds of the iteration complexity.
First, consider the case $\epsilon_{\rm sp}^{\ast}>0$. Our goal will be to choose $T$ large enough to drive the first term on the right side of (8) to below $4\frac{CL}{c\mu}\bar{\eta}\epsilon_{\rm sp}^{\ast}$. To simplify exposition, let $\kappa=\frac{CL}{c\mu}$. Then, we need to ensure that
$T\log(1-\bar{\eta}/\kappa)\leq\log\left(\frac{2\kappa\bar{\eta}\epsilon_{\rm sp}^{\ast}}{CL\|\bm{\beta}_{0}-\bm{\beta}_{\mathrm{avg}}^{\ast}\|^{2}}\right).$
@@
$T\geq\frac{\kappa}{\bar{\eta}}\log\left(\frac{CL\|\bm{\beta}_{0}-\bm{\beta}_{\mathrm{avg}}^{\ast}\|^{2}}{2\kappa\bar{\eta}\epsilon_{\rm sp}^{\ast}}\right).$
@@
$\mathbb{E}\left[\epsilon_{\rm sp}(\bm{\beta}_{T})\right]$ $\leq\ 2CL\left(1-\bar{\eta}\frac{c\mu}{CL}\right)^{T}\|\bm{\beta}_{0}-\bm{\beta}_{\mathrm{avg}}^{\ast}\|^{2}.$
@@
$T\log(1-\bar{\eta}/\kappa)\leq\log\left(\frac{\epsilon}{2CL\|\bm{\beta}_{0}-\bm{\beta}_{\mathrm{avg}}^{\ast}\|^{2}}\right).$
@@
$T\geq\frac{\kappa}{\bar{\eta}}\log\left(\frac{2CL\|\bm{\beta}_{0}-\bm{\beta}_{\mathrm{avg}}^{\ast}\|^{2}}{\epsilon}\right).$